OpenStudy (anonymous):

Fraction 4 over 5n = Fraction 2 over 3. n = ___? Fraction 2 over 15 Fraction 5 over 6 1Fraction 1 over 5 1Fraction 7 over 15

OpenStudy (anonymous):

\[\frac{ 4 }{ 5n } = \frac{ 2 }{ 3 }\] cross multiply and solve for n

OpenStudy (anonymous):

@yajna i got 8 over 15

OpenStudy (anonymous):

no wrong !! when you cross multiply, you'll get 10*n= 4*3 10n = 12 n = 12 /10 which is 1 fraction 1 over 5

OpenStudy (anonymous):

u there ?